Find the x-intercept and y-intercept of the line given by the equation.
7x-7y+6=0







Answer by mananth(16949)

7x-7y+6=0
7x-7y=-6


Find the Y & X intercepts of the two lines by plugging x=0, & y=0
equation
when x= 0,y=0.86(0, 6/7 )
when y =0,x=-0.86(-6/7 ,0)